# Discrete-modulated continuous-variable quantum key distribution secure against general attacks

### v2

@@ III Protocol
    Parameter estimation: For each $j\in[N]$, Bob announces $T_{j}$. If $T_{j}=0$, Bob announces the rounds $j$ in which $Z_{j}=\varnothing$. For each $T_{j}=1$, Alice announces $X_{j}$. Bob computes $C_{j}=f_{\mathrm{PE}}(T_{j},X_{j},Z_{j})$ for some deterministic function $f_{\mathrm{PE}}$ where $C_{j}=\perp$ if and only if $T_{j}=0$ and $C_{j}=\varnothing$ if and only if $T_{j}=1$ and $Z_{j}=\top$. If $\mathrm{freq}_{\mathbf{C}}\in\mathcal{F}_{\mathrm{tol}}$, they continue to the next step, else they abort.
@@ IV.3.3 Constructing min-tradeoff functions via dimension reduction
$g(q)=\tilde{g}(q)-g_{\mathrm{corr}}^{(\nu_{c})}(q_{\varnothing}),$ (39)
where the function $\tilde{g}(q)$ has similar properties as the function $g(q)$, except that it is evaluated on the output state of the truncated GEAT channel, where we applied the CPTNI map $\mathcal{T}_{j}$ at the end of Step 2 (instead of the actual GEAT channel). In other words, the function $\tilde{g}:\mathcal{P}_{\mathcal{C}\setminus\{\perp\}}\rightarrow\mathbb{R}$ is an affine function that gives the lower bound on the conditional von Neumann entropy evaluated on the truncated state. The function $g_{\mathrm{corr}}^{(\nu_{c})}(q_{\varnothing})$ gives the appropriate correction term and it only depends on $q_{\varnothing}$, which is the probability that $C_{j}=\varnothing$. Importantly, unlike the function $g(q)$, the function $\tilde{g}(q)$ can be computed using a finite-dimensional optimisation.
@@
$g(q)=\tilde{g}(q)-g_{\mathrm{corr}}^{(\nu_{c})}(q_{\varnothing}),$
@@
$\begin{split}\tilde{g}^{\prime}(q)\leq\inf_{\tilde{\rho}}\,&H(Z|L,E)_{\tilde{\rho}_{ZLE}}\\ \text{s.t.}\,&\tilde{\rho}_{AB}\succeq 0\\ &\mathrm{Tr}[\tilde{\rho}_{AB}]\leq 1,\\ &\mathrm{Tr}[\tilde{\rho}_{AB}]\geq 1-\kappa q_{\varnothing},\\ &\Delta(\tilde{\rho}_{A},\mathrm{Tr}_{Q}{\lvert\Psi\rangle\!\langle\Psi\rvert}_{AQ})\leq\frac{1}{2}\kappa q_{\varnothing}\\ &\mathrm{Tr}[\tilde{\rho}_{AB}\tilde{\Pi}_{c}]\leq q_{c}-\xi_{U}(q_{\varnothing})\quad\forall c\in\mathcal{C}\setminus\{\perp\},\\ &\mathrm{Tr}[\tilde{\rho}_{AB}\tilde{\Pi}_{c}]\geq q_{c}-\xi_{L}(q_{\varnothing})\quad\forall c\in\mathcal{C}\setminus\{\perp\},\end{split}$ (40)
@@
$g^{(\nu_{c})}_{\mathrm{corr}}(q_{\varnothing})$ $=m_{\mathrm{corr}}(q_{\varnothing}-\nu_{c})+c_{\mathrm{corr}},$
@@
$\xi_{U}(q_{\varnothing})$ $\geq\hat{\xi}_{U}^{(\nu_{U})}(q_{\varnothing})=\left(1-\frac{(1-2\kappa\nu_{U})}{\sqrt{\kappa\nu_{U}(1-\kappa\nu_{U})}}\right)\kappa q_{\varnothing}-\sqrt{\frac{\kappa\nu_{U}}{1-\kappa\nu_{U}}},$
$\xi_{L}(q_{\varnothing})$ $\leq\hat{\xi}_{L}^{(\nu_{L})}(q_{\varnothing})=\left(1+\frac{(1-2\kappa\nu_{L})}{\sqrt{\kappa\nu_{L}(1-\kappa\nu_{L})}}\right)\kappa q_{\varnothing}+\sqrt{\frac{\kappa\nu_{L}}{1-\kappa\nu_{L}}},$
@@ IV.3.4 SDP method to bound the conditional entropy
$\begin{split}\inf_{\sigma,\{\omega_{i,z},\eta_{i,z},\theta_{i,z}\}_{i,z},\zeta_{1},\zeta_{2}}&(1-\gamma)\sum_{i=1}^{m}\frac{w_{i}}{t_{i}\ln 2}\mathrm{Tr}\left[\sigma P_{PS}+\sum_{z\neq\varnothing}(\mathds{1}_{A}\otimes P_{z|0})(\omega_{i,z}+\omega_{i,z}^{\dagger}+(1-t_{i})\eta_{i,z})+t_{i}\theta_{i,z}P_{PS}\right]\\ \text{s.t.}\qquad&\mathrm{Tr}[\sigma]\leq 1\\ &\mathrm{Tr}[\sigma]\geq 1-\kappa q_{\varnothing}\\ &\mathrm{Tr}[\zeta_{1}+\zeta_{2}]\leq\kappa q_{\varnothing}\\ &\mathrm{Tr}[\sigma\tilde{\Pi}_{c}]\leq q_{c}-\xi_{U}(q_{\varnothing})\quad\forall c\in\mathcal{C}\setminus\{\perp\},\\ &\mathrm{Tr}[\sigma\tilde{\Pi}_{c}]\geq q_{c}-\xi_{L}(q_{\varnothing})\quad\forall c\in\mathcal{C}\setminus\{\perp\},\\ &\begin{pmatrix}\sigma&\omega_{i,z}\\ \omega_{i,z}^{\dagger}&\eta_{i,z}\end{pmatrix}\succeq 0\quad\forall i,z\\ &\begin{pmatrix}\sigma&\omega_{i,z}^{\dagger}\\ \omega_{i,z}&\theta_{i,z}\end{pmatrix}\succeq 0\quad\forall i,z\\ &\begin{pmatrix}\zeta_{1}&\mathrm{Tr}_{B}[\sigma]-\mathrm{Tr}_{Q}[{\lvert\Psi\rangle\!\langle\Psi\rvert}]\\ \mathrm{Tr}_{B}[\sigma]-\mathrm{Tr}_{Q}[{\lvert\Psi\rangle\!\langle\Psi\rvert}]&\zeta_{2}\end{pmatrix}\succeq 0\end{split}$ (48)
@@ IV.3.5 Dual solution
$\lambda_{\text{norm}}:\quad$ $1-\mathrm{Tr}[\sigma]\leq\kappa q_{\varnothing}$
$\lambda_{\text{dist}}:\quad$ $\mathrm{Tr}[\zeta_{1}+\zeta_{2}]\leq\kappa q_{\varnothing}$
$\lambda_{c}^{U}:\quad$ $\mathrm{Tr}[\sigma\tilde{\Pi}_{c}]\leq q_{c}-\xi_{U}(q_{\varnothing})$
$\lambda^{L}_{c}:\quad$ $\mathrm{Tr}[\sigma\tilde{\Pi}_{c}]\geq q_{c}-\xi_{L}(q_{\varnothing})$
@@
$\tilde{g}^{\prime}(q)\geq\varphi-\lambda_{\text{norm}}\kappa q_{\varnothing}-\lambda_{\text{dist}}\kappa q_{\varnothing}\\ +\sum_{c\neq\perp,\varnothing}\left(-\lambda_{c}^{U}(q_{c}-\xi_{U}(q_{\varnothing}))+\lambda_{c}^{L}(q_{c}-\xi_{L}(q_{\varnothing}))\right)\\ -\lambda_{\varnothing}^{U}(q_{\varnothing}-\xi_{U}(q_{\varnothing}))+\lambda_{\varnothing}^{L}(q_{\varnothing}-\xi_{L}(q_{\varnothing})),$ (49)
@@
$\begin{split}\xi_{U}(q_{\varnothing})&\geq m_{U}q_{\varnothing}+c_{U}\\ \xi_{L}(q_{\varnothing})&\leq m_{L}q_{\varnothing}+c_{L},\end{split}$ (50)
@@
$\tilde{g}(q)\geq\varphi^{\prime}+\sum_{c\neq\perp,\varnothing}(-\lambda^{U}_{c}+\lambda_{c}^{L})q_{c}\\ +q_{\varnothing}\Bigg{(}-\kappa\lambda_{\text{norm}}-\kappa\lambda_{\text{dist}}-\lambda^{U}_{\varnothing}+\lambda^{L}_{\varnothing}\\ +\sum_{c\neq\perp}(\lambda_{c}^{U}m_{U}-\lambda_{c}^{L}m_{L})\Bigg{)},$ (51)
@@
$\lambda^{\prime}_{c}=\begin{cases}\lambda_{c}^{L}-\lambda_{c}^{U}&\text{if $c\neq\perp,\varnothing$},\\ -\lambda_{\varnothing}^{U}+\lambda_{\varnothing}^{L}-\kappa\lambda_{\text{norm}}-\kappa\lambda_{\text{dist}}+\sum_{c\neq\perp}(-\lambda_{c}^{L}m_{L}+\lambda_{c}^{U}m_{U})-m_{\mathrm{corr}}&\text{if $c=\varnothing$}.\end{cases}$ (54)
@@ Acknowledgements.
Wen Yu Kon and Charles Lim, in their present role at JPMorgan Chase & Co contributed to this work for information purposes. This paper is not a product of the Research Department of J.P. Morgan, Singapore, or its affiliates. Neither J.P. Morgan, Singapore nor any of its affiliates make any explicit or implied representation or warranty and none of them accept any liability in connection with this paper, including, but limited to, the completeness, accuracy, reliability of information contained herein and the potential legal, compliance, tax or accounting effects thereof. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction.
@@ A.2 Bounding the trace distance between truncated and original state
where the last term can be monitored in the test rounds as $\mathrm{freq}(C=\varnothing)$. Let us introduce the shorthand $\Pr[W>\tau_{\mathrm{max}}]=\nu$, and $\kappa=\Gamma(n_{\mathrm{max}}+2,0)/\Gamma(n_{\mathrm{max}}+2,\tau_{\mathrm{max}})$. We can write $\delta=\sqrt{2\kappa\nu}$.
@@ A.5 Proof of dimension reduction theorem
$\begin{split}\tilde{g}^{\prime}(q)=\inf_{\sigma_{AB}}\quad&H(Z|L,E)_{\mathcal{N}[\sigma_{AB}]}\\ \text{s.t.}\quad&\sigma_{AB}\succeq 0,\\ &\mathrm{Tr}[\sigma_{AB}]\leq 1,\\ &\mathrm{Tr}[\sigma_{AB}]\geq 1-\kappa q_{\varnothing},\\ &\Delta(\sigma_{A},\mathrm{Tr}_{Q}[{\lvert\Psi\rangle\!\langle\Psi\rvert}_{AQ}])\leq\kappa q_{\varnothing}/2,\\ &\mathrm{Tr}[\sigma_{AB}\Pi_{c}]\leq q_{c}-\xi_{U}(q_{\varnothing}),\\ &\mathrm{Tr}[\sigma_{AB}\Pi_{c}]\geq q_{c}+\xi_{L}(q_{\varnothing})\end{split}$ (106)
@@
$\geq\inf_{\sigma_{AB}:\,\Delta(\rho_{AB},\sigma_{AB})\leq\delta(q_{\varnothing}),\,\rho_{AB}\in\mathcal{S}_{q}}H(Z|L,E)_{\mathcal{N}[\sigma_{AB}]}\underbrace{-\delta(q_{\varnothing})\log_{2}d_{Z}-(1+\delta(q_{\varnothing}))h_{2}\left(\frac{\delta(q_{\varnothing})}{1+\delta(q_{\varnothing})}\right)}_{-v(\delta)}$
$\geq\tilde{g}(q)-g_{\mathrm{corr}}^{(\nu_{c})}(q_{\varnothing}).$ (107)
Here, we denote $\mathcal{S}_{q}$ as a shorthand for the feasible set defined in Eq. (64). The second line is due to Lemma 2. The last line can be deduced by identifying that the constraints $\Delta(\rho_{AB},\sigma_{AB})\leq\delta(q_{\varnothing})$ and $\rho_{AB}\in\mathcal{S}_{q}$ are equivalent to the one written in Eq. (106). Additionally, we notice that the correction terms $v(\delta(q_{\varnothing}))$ is upper bounded by $g_{\mathrm{corr}}^{(\nu_{c})}(q_{\varnothing})$, as derived previously. This concludes the proof of Theorem 2.

### v3

@@ III Protocol
    Parameter estimation: For each $j\in[N]$, Bob announces $T_{j}$. If $T_{j}=0$, Bob announces the rounds $j$ in which $Z_{j}=\varnothing$. For each $T_{j}=1$, Alice announces $X_{j}$. Bob computes $C_{j}=f_{\mathrm{PE}}(T_{j},X_{j},Z_{j})$ for some deterministic function $f_{\mathrm{PE}}$ where $C_{j}=\perp$ if and only if $T_{j}=0$ and $C_{j}=\top$ if and only if $T_{j}=1$ and $Z_{j}=\top$. If $\mathrm{freq}_{\mathbf{C}}\in\mathcal{F}_{\mathrm{tol}}$, they continue to the next step, else they abort.
@@ IV.3.3 Constructing min-tradeoff functions via dimension reduction
$g(q)=\tilde{g}(q)-g_{\mathrm{corr}}^{(\nu_{c})}(q_{\top}),$ (39)
where the function $\tilde{g}(q)$ has similar properties as the function $g(q)$, except that it is evaluated on the output state of the truncated GEAT channel, where we applied the CPTNI map $\mathcal{T}_{j}$ at the end of Step 2 (instead of the actual GEAT channel). In other words, the function $\tilde{g}:\mathcal{P}_{\mathcal{C}\setminus\{\perp\}}\rightarrow\mathbb{R}$ is an affine function that gives the lower bound on the conditional von Neumann entropy evaluated on the truncated state. The function $g_{\mathrm{corr}}^{(\nu_{c})}(q_{\top})$ gives the appropriate correction term and it only depends on $q_{\top}$, which is the probability that $C_{j}=\top$. Importantly, unlike the function $g(q)$, the function $\tilde{g}(q)$ can be computed using a finite-dimensional optimisation.
@@
$g(q)=\tilde{g}(q)-g_{\mathrm{corr}}^{(\nu_{c})}(q_{\top}),$
@@
$\begin{split}\tilde{g}^{\prime}(q)\leq\inf_{\tilde{\rho}}\,&H(Z|L,E)_{\tilde{\rho}_{ZLE}}\\ \text{s.t.}\,&\tilde{\rho}_{AB}\succeq 0\\ &\mathrm{Tr}[\tilde{\rho}_{AB}]\leq 1,\\ &\mathrm{Tr}[\tilde{\rho}_{AB}]\geq 1-\kappa q_{\top},\\ &\Delta(\tilde{\rho}_{A},\mathrm{Tr}_{Q}{\lvert\Psi\rangle\!\langle\Psi\rvert}_{AQ})\leq\frac{1}{2}\kappa q_{\top}\\ &\mathrm{Tr}[\tilde{\rho}_{AB}\tilde{\Pi}_{c}]\leq q_{c}-\xi_{U}(q_{\top})\quad\forall c\in\mathcal{C}\setminus\{\perp\},\\ &\mathrm{Tr}[\tilde{\rho}_{AB}\tilde{\Pi}_{c}]\geq q_{c}-\xi_{L}(q_{\top})\quad\forall c\in\mathcal{C}\setminus\{\perp\},\end{split}$ (40)
@@
$g^{(\nu_{c})}_{\mathrm{corr}}(q_{\top})$ $=m_{\mathrm{corr}}(q_{\top}-\nu_{c})+c_{\mathrm{corr}},$
@@
$\xi_{U}(q_{\top})$ $\geq\hat{\xi}_{U}^{(\nu_{U})}(q_{\top})=\left(1-\frac{(1-2\kappa\nu_{U})}{\sqrt{\kappa\nu_{U}(1-\kappa\nu_{U})}}\right)\kappa q_{\top}-\sqrt{\frac{\kappa\nu_{U}}{1-\kappa\nu_{U}}},$
$\xi_{L}(q_{\top})$ $\leq\hat{\xi}_{L}^{(\nu_{L})}(q_{\top})=\left(1+\frac{(1-2\kappa\nu_{L})}{\sqrt{\kappa\nu_{L}(1-\kappa\nu_{L})}}\right)\kappa q_{\top}+\sqrt{\frac{\kappa\nu_{L}}{1-\kappa\nu_{L}}},$
@@ IV.3.4 SDP method to bound the conditional entropy
$\begin{split}\inf_{\sigma,\{\omega_{i,z},\eta_{i,z},\theta_{i,z}\}_{i,z},\zeta_{1},\zeta_{2}}&(1-\gamma)\sum_{i=1}^{m}\frac{w_{i}}{t_{i}\ln 2}\mathrm{Tr}\left[\sigma P_{PS}+\sum_{z\neq\varnothing}(\mathds{1}_{A}\otimes P_{z|0})(\omega_{i,z}+\omega_{i,z}^{\dagger}+(1-t_{i})\eta_{i,z})+t_{i}\theta_{i,z}P_{PS}\right]\\ \text{s.t.}\qquad&\mathrm{Tr}[\sigma]\leq 1\\ &\mathrm{Tr}[\sigma]\geq 1-\kappa q_{\top}\\ &\mathrm{Tr}[\zeta_{1}+\zeta_{2}]\leq\kappa q_{\top}\\ &\mathrm{Tr}[\sigma\tilde{\Pi}_{c}]\leq q_{c}-\xi_{U}(q_{\top})\quad\forall c\in\mathcal{C}\setminus\{\perp\},\\ &\mathrm{Tr}[\sigma\tilde{\Pi}_{c}]\geq q_{c}-\xi_{L}(q_{\top})\quad\forall c\in\mathcal{C}\setminus\{\perp\},\\ &\begin{pmatrix}\sigma&\omega_{i,z}\\ \omega_{i,z}^{\dagger}&\eta_{i,z}\end{pmatrix}\succeq 0\quad\forall i,z\\ &\begin{pmatrix}\sigma&\omega_{i,z}^{\dagger}\\ \omega_{i,z}&\theta_{i,z}\end{pmatrix}\succeq 0\quad\forall i,z\\ &\begin{pmatrix}\zeta_{1}&\mathrm{Tr}_{B}[\sigma]-\mathrm{Tr}_{Q}[{\lvert\Psi\rangle\!\langle\Psi\rvert}]\\ \mathrm{Tr}_{B}[\sigma]-\mathrm{Tr}_{Q}[{\lvert\Psi\rangle\!\langle\Psi\rvert}]&\zeta_{2}\end{pmatrix}\succeq 0\end{split}$ (48)
@@ IV.3.5 Dual solution
$\lambda_{\text{norm}}:\quad$ $1-\mathrm{Tr}[\sigma]\leq\kappa q_{\top}$
$\lambda_{\text{dist}}:\quad$ $\mathrm{Tr}[\zeta_{1}+\zeta_{2}]\leq\kappa q_{\top}$
$\lambda_{c}^{U}:\quad$ $\mathrm{Tr}[\sigma\tilde{\Pi}_{c}]\leq q_{c}-\xi_{U}(q_{\top})$
$\lambda^{L}_{c}:\quad$ $\mathrm{Tr}[\sigma\tilde{\Pi}_{c}]\geq q_{c}-\xi_{L}(q_{\top})$
@@
$\tilde{g}^{\prime}(q)\geq\varphi-\lambda_{\text{norm}}\kappa q_{\top}-\lambda_{\text{dist}}\kappa q_{\top}\\ +\sum_{c\neq\perp,\top}\left(-\lambda_{c}^{U}(q_{c}-\xi_{U}(q_{\top}))+\lambda_{c}^{L}(q_{c}-\xi_{L}(q_{\top}))\right)\\ -\lambda_{\top}^{U}(q_{\top}-\xi_{U}(q_{\top}))+\lambda_{\top}^{L}(q_{\top}-\xi_{L}(q_{\top})),$ (49)
@@
$\begin{split}\xi_{U}(q_{\top})&\geq m_{U}q_{\top}+c_{U}\\ \xi_{L}(q_{\top})&\leq m_{L}q_{\top}+c_{L},\end{split}$ (50)
@@
$\tilde{g}(q)\geq\varphi^{\prime}+\sum_{c\neq\perp,\top}(-\lambda^{U}_{c}+\lambda_{c}^{L})q_{c}\\ +q_{\top}\Bigg{(}-\kappa\lambda_{\text{norm}}-\kappa\lambda_{\text{dist}}-\lambda^{U}_{\top}+\lambda^{L}_{\top}\\ +\sum_{c\neq\perp}(\lambda_{c}^{U}m_{U}-\lambda_{c}^{L}m_{L})\Bigg{)},$ (51)
@@
$\lambda^{\prime}_{c}=\begin{cases}\lambda_{c}^{L}-\lambda_{c}^{U}&\text{if $c\neq\perp,\top$},\\ -\lambda_{\top}^{U}+\lambda_{\top}^{L}-\kappa\lambda_{\text{norm}}-\kappa\lambda_{\text{dist}}+\sum_{c\neq\perp}(-\lambda_{c}^{L}m_{L}+\lambda_{c}^{U}m_{U})-m_{\mathrm{corr}}&\text{if $c=\top$}.\end{cases}$ (54)
@@ VI Discussion and conclusion
It is also possible that this variant of DM-CV-QKD protocol has significant finite-size penalties when analysed using GEAT. This may be the consequence of the linearisation terms. In general, the probability of exceeding $\tau_{\max}$ is small, indicating a low probability of having large energy. As such, the choice of linearisation parameters $\nu_{U}$ and $\nu_{L}$ would be correspondingly small as well to have a small gap between the actual value (e.g. $\xi_{L}$) and linearised value (e.g. $\hat{\xi}_{L}^{(\nu_{L})}$). Since this term is present in the denominator of the gradient, it leads to a large gradient and causes the penalty from GEAT to be significant. As the finite-size penalty in both EAT and GEAT are very sensitive to the gradient of the min-tradeoff function, we expect that the huge finite-size penalty is a feature of the DM-CV-QKD protocol which estimates the dimension reduction correction using the probability of exceeding $\tau_{\max}$. Therefore, to improve the finite-size penalty, we may require a new technique that is less sensitive to the gradient of the min-tradeoff function. Alternatively, we may require another method to estimate the dimension reduction corrections. A possible approach is to investigate continuity bounds that are based on fidelities instead of trace distances. For example, in the case for quantum-classical states (instead of the case for classical-quantum states that we need here), it is known that the continuity bound based on fidelity [51] has a better correction term. Another potential room for improvement is to find corrections to the statistical constraints that have gentler gradients at $q_{\top}\approx 0.$ Recently, Ref. [52] proposed the Rényi entropy version of the GEAT. Interestingly, this version of the entropy accumulation theorem does not require the usual analysis via affine min-tradeoff functions, which removes the necessity of linearising the correction terms. This approach also removes the need to find the best $q_{\mathrm{dual}}$, which might potentially reduce the minimum block size $N$.
@@ Acknowledgements.
The authors acknowledge funding support from the National Research Foundation of Singapore (NRF) Fellowship grant (NRFF11-2019-0001) and NRF Quantum Engineering Programme 1.0 grant (QEP-P2 and QEP-P3). We also thank Ernest Y.-Z. Tan for discussion and his valuable feedback on an earlier version of this manuscript. Wen Yu Kon and Charles Lim, in their present role at JPMorgan Chase & Co contributed to this work for information purposes. This paper is not a product of the Research Department of J.P. Morgan, Singapore, or its affiliates. Neither J.P. Morgan, Singapore nor any of its affiliates make any explicit or implied representation or warranty and none of them accept any liability in connection with this paper, including, but limited to, the completeness, accuracy, reliability of information contained herein and the potential legal, compliance, tax or accounting effects thereof. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction.
@@ A.2 Bounding the trace distance between truncated and original state
where the last term can be monitored in the test rounds as $\mathrm{freq}(C=\top)$. Let us introduce the shorthand $\Pr[W>\tau_{\mathrm{max}}]=\nu$, and $\kappa=\Gamma(n_{\mathrm{max}}+2,0)/\Gamma(n_{\mathrm{max}}+2,\tau_{\mathrm{max}})$. We can write $\delta=\sqrt{2\kappa\nu}$.
@@ A.5 Proof of dimension reduction theorem
$\begin{split}\tilde{g}^{\prime}(q)=\inf_{\sigma_{AB}}\quad&H(Z|L,E)_{\mathcal{N}[\sigma_{AB}]}\\ \text{s.t.}\quad&\sigma_{AB}\succeq 0,\\ &\mathrm{Tr}[\sigma_{AB}]\leq 1,\\ &\mathrm{Tr}[\sigma_{AB}]\geq 1-\kappa q_{\top},\\ &\Delta(\sigma_{A},\mathrm{Tr}_{Q}[{\lvert\Psi\rangle\!\langle\Psi\rvert}_{AQ}])\leq\kappa q_{\top}/2,\\ &\mathrm{Tr}[\sigma_{AB}\Pi_{c}]\leq q_{c}-\xi_{U}(q_{\top}),\\ &\mathrm{Tr}[\sigma_{AB}\Pi_{c}]\geq q_{c}+\xi_{L}(q_{\top})\end{split}$ (106)
@@
$\geq\inf_{\sigma_{AB}:\,\Delta(\rho_{AB},\sigma_{AB})\leq\delta(q_{\top}),\,\rho_{AB}\in\mathcal{S}_{q}}H(Z|L,E)_{\mathcal{N}[\sigma_{AB}]}\underbrace{-\delta(q_{\top})\log_{2}d_{Z}-(1+\delta(q_{\top}))h_{2}\left(\frac{\delta(q_{\top})}{1+\delta(q_{\top})}\right)}_{-v(\delta)}$
$\geq\tilde{g}(q)-g_{\mathrm{corr}}^{(\nu_{c})}(q_{\top}).$ (107)
Here, we denote $\mathcal{S}_{q}$ as a shorthand for the feasible set defined in Eq. (64). The second line is due to Lemma 2. The last line can be deduced by identifying that the constraints $\Delta(\rho_{AB},\sigma_{AB})\leq\delta(q_{\top})$ and $\rho_{AB}\in\mathcal{S}_{q}$ are equivalent to the one written in Eq. (106). Additionally, we notice that the correction terms $v(\delta(q_{\top}))$ is upper bounded by $g_{\mathrm{corr}}^{(\nu_{c})}(q_{\top})$, as derived previously. This concludes the proof of Theorem 2.